5. The ruler is in equilibrium. The hanging load on the left arm of the visor, at point A1, has mass m = 500 g. What force does the force gauge hook apply to point A2 of the ruler? It is known that d1 = 10 cm, while d2 = 1/2 × d1


1
Expert's answer
2021-02-22T10:26:44-0500
"F_1d_1=F_2d_2,""m_1gd_1=F_2\dfrac{1}{2}d_1,""F_2=2m_1g=2\cdot0.5\ kg\cdot9.8\ \dfrac{m}{s^2}=9.8\ N."